Today's maximum wind gust was 8 mph less than 4 times the average wind speed. If today's maximum wind gust was twice yesterday's maximum of 32 mph, what was today's average wind speed?

Answer by ankor@dixie-net.com(22740)

Today's maximum wind gust was 8 mph less than 4 times the average wind speed.
M = 4A - 8
If today's maximum wind gust was twice yesterday's maximum of 32 mph, what was today's average wind speed?
M = 2(32)
M = 64
4A - 8 = 64
4A = 64 + 8
A = 72/4
A = 18 mph is todays av wind